OpenStudy (anonymous):

What is the sum of three and three over eight plus two and three over four in simplest form? six and one over eight five and one over two five and three over four six and one over four

OpenStudy (anonymous):

3 3/8+2 3/4 this is the equation right?

OpenStudy (anonymous):

l= yes it is

OpenStudy (anonymous):

k well we need the denominator for both of them to be the same

OpenStudy (anonymous):

srry dont mind the equal sign my backsdpace dont work

OpenStudy (anonymous):

what is the closest denominator they both have?

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

look lemme draw a pic to make it a little simpler

OpenStudy (anonymous):



OpenStudy (anonymous):

so its 8 so to get the right side denominator to 8 multiply the fraction by 2.

OpenStudy (anonymous):

which would be 3 3/8+ 2 6/8

OpenStudy (anonymous):

that would equal 5 9/8 now we simplyfy it which would be 5 1/8

OpenStudy (anonymous):

hope this helped :3

OpenStudy (anonymous):

@keke12147

OpenStudy (anonymous):

@chucho78 \[5+\frac{9}{8}=5+1+\frac{1}{8}=6+\frac{1}{8} \]

OpenStudy (anonymous):

yeah sorry! it i forgot to add the one! @robtobey